OpenStudy (anonymous):

Choose the procedure you would use to solve the equation 5x + 3 = 28 for x

OpenStudy (lilg132):

5x + 3 = 28 5x = 28 - 3 5x = 25 5x / 5 = 25 / 5 x = 5 remember to click good answer

jhonyy9 (jhonyy9):

5x=25 x=5